### 17 Mar 2026 v1

# Semiparametric Inference under Dual Positivity Boundaries: Nested Identification with Administrative Censoring and Confounded Treatment

Abstract
When a long-term outcome is administratively censored for a substantial fraction of a study cohort while a short-term intermediate variable remains broadly available, the target causal parameter can be identified through a nested functional that integrates the outcome regression over the conditional intermediate distribution, avoiding inverse censoring weights entirely. In observational studies where treatment is also confounded, this nested identification creates a semiparametric structure with two distinct positivity boundaries — one from the censoring mechanism and one from the treatment assignment — that enter the efficient influence function in fundamentally different roles. The censoring boundary is removed from the identification by the nested functional but remains in the efficient score; the treatment boundary appears in both.
We develop the inference theory for this dual-boundary structure. Three results are established. First, a fluctuation coupling proposition shows that when a targeted estimator solves the nested score equation, the rate of the intermediate regression depends on the outcome model quality but not on the treatment propensity quality at leading order, because the score residuals have conditional mean zero under any propensity weighting (Proposition 1). Second, the doubly robust remainder decomposes into two bilinear products with an asymmetric three-component robustness geometry: the treatment propensity must be consistent at any rate for $\sqrt{n}$-consistency, but correct outcome modelling alone — sufficient in the single-boundary case — places the estimator at the product-rate boundary when the propensity is misspecified, producing a near-boundary regime with persistent sandwich undercoverage (Theorem 1). Third, the two remainder components can interact through a covariance term $c_{12}$ that is positive when both boundaries concentrate in the same covariate region, amplifying the variance gap beyond the sum of individual contributions (Proposition 2). The leave-one-out jackknife captures the full dual-boundary variance and restores nominal coverage uniformly across all regimes. Simulations calibrated to observational comparative effectiveness studies confirm that the dual-boundary regime produces variance inflation exceeding the single-boundary case, and that the nested identification strategy achieves lower near-boundary severity than standard inverse-weighted estimators when administrative censoring is the dominant positivity concern.
Keywords: semiparametric inference; positivity boundary; nested identification; administrative censoring; doubly robust estimation; variance decomposition; jackknife; targeted learning.
1 Introduction
The missing-data structure
A recurring structure in longitudinal studies and comparative effectiveness research involves two distinct sources of incomplete observation operating simultaneously. The first is administrative censoring: a primary outcome requires extended follow-up, and a non-trivial fraction of individuals exit the study before the outcome is ascertained, creating a censoring probability $g_{\Delta}$ that can approach zero for late enrollees or short-enrollment subgroups. The second is treatment confounding: in observational settings, treatment assignment depends on measured covariates, and the propensity score $g_{A}$ can approach zero or one in subgroups with strongly confounded treatment selection.
Each source alone is well understood. Administrative censoring is addressed by inverse probability of censoring weighting (IPCW; Robins et al. 1994; Robins and Rotnitzky 1995), which becomes unstable as $g_{\Delta}\to 0$. Treatment confounding is addressed by inverse propensity weighting or doubly robust augmentation (Bang and Robins 2005; Chernozhukov et al. 2018), which becomes unstable as $g_{A}\to 0$ or $1$. Standard doubly robust estimators combine both corrections by inverse-weighting with the product $g_{\Delta}^{-1}g_{A}^{-1}$, compounding the two positivity concerns.
The interaction between these two sources — when they are simultaneously active and potentially concentrated in the same covariate region — has received less theoretical attention. This paper develops the inference theory for that interaction in the context of a nested identification strategy that separates the two boundaries into distinct roles.
Prevalence
The dual-boundary structure is common. In pharmacoepidemiologic studies of statin initiation, nearly half of elderly patients discontinue therapy within the first year (Benner et al., 2002); those who disenroll contribute no long-term cardiovascular outcome data while their 3-month LDL cholesterol is already recorded. In comparative effectiveness studies of diabetes therapies, Dagan et al. (2018) report that over 58% of a 40,000-patient cohort were censored within 12 months. In both settings, a short-term biomarker is available for a substantially larger fraction of the cohort because it is measured before the censoring event occurs.
Nested identification and the two boundaries
When a short-term intermediate variable $S$ satisfies a surrogate-mediated missing-at-random condition ($\Delta\perp Y\mid S,A,W$), the causal parameter can be identified through a nested functional:
$\Psi(P_{0})=E_{W}\bigl[E_{S|A=1,W}[\bar{Q}_{Y}(S,1,W)]-E_{S|A=0,W}[\bar{Q}_{Y}(S,0,W)]\bigr],$ (1)
where $\bar{Q}_{Y}(S,a,W)=E[Y\mid S,A\!=\!a,W,\Delta\!=\!1]$ (Gilbert et al. 2008; Athey et al. 2020; Li 2026a). This avoids $g_{\Delta}^{-1}$ entirely: the censoring propensity does not appear in the identification.
The treatment propensity $g_{A}$, however, remains in the efficient influence function through the clever covariates that define the targeting step. This creates a structural asymmetry: one boundary (censoring) is removed from the identification but retained in the score; the other (treatment) appears in both. The resulting structure differs from both the classical single-propensity doubly robust problem and from the multiply robust framework of Tchetgen Tchetgen and Shpitser (2012), where nuisance components enter symmetrically.
A further complication is specific to nested functionals. The identification (1) involves an outer integration over the conditional distribution of $S$. A one-step debiased estimator plugging in separate estimates $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ produces a cross-product remainder $R_{SY}$ with no doubly robust complement that cross-fitting does not remove (Li 2026a, Proposition 1). A two-stage targeting construction can absorb $R_{SY}$ algebraically without estimating $f_{S}$, but in the observational setting the targeting step uses $\hat{g}_{A}$-weighted empirical means, coupling the intermediate regression to the propensity quality.
Contributions
We make four contributions to semiparametric missing-data methodology.
First, the fluctuation coupling (Proposition 3). When a targeted estimator solves the nested score equation, the post-targeting intermediate regression converges at rate $O_{P}(r_{Q}+n^{-1/2})$ regardless of propensity quality. The mechanism is that the score residuals have conditional mean zero given the variables that the propensity depends on, so any propensity reweighting preserves the CLT rate. The actual cost of wrong propensity weights is not bias but an irreducible $O_{P}(n^{-1/2})$ term that places the second remainder at the product-rate boundary.
Second, the three-component robustness geometry (Theorem 4, Corollary 5). The remainder decomposes into $R_{2}^{(1)}$ (outcome $\times$ censoring) and $R_{2}^{(2)}$ (intermediate $\times$ treatment). The treatment propensity must be consistent at any rate — a necessary condition with no analog in the single-boundary case. The geometry is three-component but not multiply robust in the sense of Tchetgen Tchetgen and Shpitser (2012): no single correct component guarantees $\sqrt{n}$-consistency.
Third, the dual near-boundary variance decomposition (Proposition 6). The sampling variance includes two remainder-variance terms and their covariance $c_{12}$. We show $c_{12}>0$ when both boundaries concentrate in the same covariate region, so the total gap exceeds the sum of individual contributions.
Fourth, uniform jackknife validity (Theorem 10). The jackknife is consistent for the total variance across all four regimes (strong decay, single-censoring, single-treatment, dual) and the Wald interval achieves exact coverage uniformly.
Related work
Doubly robust estimation. The classical theory (Robins et al. 1994; Bang and Robins 2005; Chernozhukov et al. 2018) addresses a single missing-data mechanism with two-component robustness. The nested identification generates a remainder $R_{SY}$ with no doubly robust complement, and the observational treatment creates a three-component structure.
Multiply robust estimation. Tchetgen Tchetgen and Shpitser (2012) develop estimators consistent if any one of $K$ models is correct. The geometry here is different: no single component suffices. This is a form of asymmetric partial robustness.
Positivity violations. Weight truncation (Crump et al., 2009), modified treatment rules (Díaz and van der Laan, 2012), and overlap weighting (Li et al., 2018) address single-propensity boundaries. The present work extends the variance-decomposition approach of Li (2026b) from one boundary to two interacting boundaries.
Nested identification. Athey et al. (2020) develop the surrogate index without censoring or positivity concerns; Li (2026a) develop the estimator for randomized trials with known treatment assignment. Neither addresses the dual-boundary structure.
Outline
Section 2 defines the data structure and identification. Section 3 develops the semiparametric theory. Section 4 constructs the targeted estimator. Section 5 establishes the asymptotic theory. Section 6 reports simulations. Section 7 discusses limitations and extensions.
2 Nested Identification under Two Positivity Boundaries
We formalize the semiparametric structure described in Section 1. The observational cohort is the motivating instance; the identification theorem and the dual-boundary framework apply to any setting where a nested functional is combined with administrative censoring and confounded treatment.
2.1 Observed-data structure
Consider a cohort of $J$ clusters (e.g., clinics), each contributing $m$ individuals, for $N=Jm$ total observations. The observed data for individual $i$ in cluster $j$ are
$O_{ij}=(W_{ij},A_{ij},S_{ij},\Delta_{ij},\Delta_{ij}Y_{ij})\sim P_{0},$ (2)
where $W_{ij}\in\mathbb{R}^{p}$ are baseline covariates, $A_{ij}\in\{0,1\}$ is the treatment received, $S_{ij}$ is the short-term intermediate variable, $\Delta_{ij}\in\{0,1\}$ indicates whether the long-term outcome was observed, and $Y_{ij}$ is that outcome. A cluster random effect $b_{j}\sim N(0,\sigma_{b}^{2})$ induces within-cluster correlation. Cluster-level vectors $O_{j}=\{O_{ij}\}_{i=1}^{m}$ are independent across $j$.
The target parameter is the average treatment effect $\Psi(P_{0})=E[Y_{ij}(1)-Y_{ij}(0)]$.
2.2 Conditions defining the dual-boundary structure
Five conditions formalize the setting.
Assumption 1 (Consistency).
For $A_{ij}=a$: $S_{ij}=S_{ij}(a)$ and $Y_{ij}=Y_{ij}(a)$.
Assumption 2 (No unmeasured confounders).
$(Y(0),Y(1),S(0),S(1))\perp A_{ij}\mid W_{ij}$.
Assumption 3 (Surrogate-mediated MAR).
$\Delta_{ij}\perp Y_{ij}\mid S_{ij},A_{ij},W_{ij}$. This is the structural condition enabling the nested identification: once the intermediate variable is observed, censoring carries no additional information about the outcome.
Assumption 4 (Censoring support positivity).
$P(\Delta=1\mid S,A,W)\geq c_{\Delta}>0$ on the relevant support. This is a support condition for estimating $\bar{Q}_{Y}$, not an inverse-weighting condition: the nested identification avoids $g_{\Delta}^{-1}$.
Assumption 5 (Treatment positivity).
$c_{A}\leq g_{A}^{0}(a\mid w)\leq 1-c_{A}$ for some $c_{A}>0$. Unlike Assumption 4, this is an inverse-weighting condition: $g_{A}^{-1}$ appears in the targeting step. This is the second positivity boundary.
The asymmetry between Assumptions 4 and 5 is the structural source of the dual-boundary phenomenon: the nested identification removes $g_{\Delta}^{-1}$ from the target parameter (Assumption 4 is a support condition only) but retains $g_{A}^{-1}$ in the efficient score (Assumption 5 is an inverse-weighting condition).
2.3 Identification
Theorem 1 (Identification via the nested functional).
Under Assumptions 15:
$\Psi(P_{0})=E_{W}\bigl[E_{S|A=1,W}[\bar{Q}_{Y}(S,1,W)]-E_{S|A=0,W}[\bar{Q}_{Y}(S,0,W)]\bigr],$ (3)
where $\bar{Q}_{Y}(S,a,W)=E[Y\mid S,A\!=\!a,W,\Delta\!=\!1]$.
Proof.
By Assumptions 12, $E[Y(a)\mid S(a),W]=E[Y\mid S,A\!=\!a,W]$. By Assumptions 34, $E[Y\mid S,A\!=\!a,W]=E[Y\mid S,A\!=\!a,W,\Delta\!=\!1]$. Integrating over $P(S\mid a,W)$ and marginalizing gives (3). ∎
Remark 1 (What the nested functional separates).
The identification (3) does not contain $g_{\Delta}^{-1}$. The censoring propensity enters the EIF (Section 3) but not the identification. The treatment propensity $g_{A}^{-1}$ appears in the targeting clever covariates. The two boundaries therefore serve different structural roles: $g_{\Delta}$ is a nuisance for efficient estimation; $g_{A}$ is a nuisance for both identification (through Assumption 2) and estimation. This asymmetry propagates through the remainder structure and the robustness geometry.
3 Semiparametric Theory
We develop the EIF, the remainder decomposition, and the robustness structure for nested functionals under the dual-boundary conditions of Section 2.
3.1 Efficient influence function
The individual-level EIF components are, for $a\in\{0,1\}$:
$D_{Y,a}(O)$ $=\frac{\mathbb{I}(A\!=\!a)\,\Delta}{g_{A}(a\mid W)\,g_{\Delta}(S,a,W)}\bigl\{Y-\bar{Q}_{Y}(S,a,W)\bigr\},$ (4)
$D_{S,a}(O)$ $=\frac{\mathbb{I}(A\!=\!a)}{g_{A}(a\mid W)}\bigl\{\bar{Q}_{Y}(S,a,W)-\bar{Q}_{\mathrm{int}}(a,W)\bigr\},$ (5)
$D_{W,a}(O)$ $=\bar{Q}_{\mathrm{int}}(a,W)-\Psi_{a},$ (6)
with $D^{*}(O)=\sum_{a}(-1)^{1-a}(D_{Y,a}+D_{S,a}+D_{W,a})$. The derivation follows the tangent-space factorization $\mathcal{T}=\mathcal{T}_{W}\oplus\mathcal{T}_{A}\oplus\mathcal{T}_{S}\oplus\mathcal{T}_{\Delta}\oplus\mathcal{T}_{Y}$. The pathwise derivative in the $\mathcal{T}_{\Delta}$ direction vanishes under Assumption 3: the EIF has no censoring-mechanism component. In clustered data, the cluster-level EIC is $\mathrm{EIF}_{j}=\sum_{i}D^{*}(O_{ij})$.
Both boundaries are visible in the EIF: $g_{\Delta}$ enters the denominator of $D_{Y,a}$ (first boundary); $g_{A}$ enters the denominator of both $D_{Y,a}$ and $D_{S,a}$ (second boundary).
3.2 Remainder decomposition
After two-stage targeting (Section 4), the nested cross-product $R_{SY}$ involving the conditional intermediate distribution is eliminated algebraically (Lemma 2). The surviving bilinear remainder is
$R_{2}=\underbrace{\int\bigl(\hat{\bar{Q}}_{Y}^{*}-\bar{Q}_{Y}^{0}\bigr)\Bigl(\frac{\hat{g}_{\Delta}}{g_{\Delta}^{0}}-1\Bigr)\,dP_{0}}_{\displaystyle R_{2}^{(1)}:\;\text{censoring boundary}}\;+\;\underbrace{\int\bigl(\hat{\bar{Q}}_{\mathrm{int}}^{*}-\bar{Q}_{\mathrm{int}}^{0}\bigr)\Bigl(\frac{\hat{g}_{A}}{g_{A}^{0}}-1\Bigr)\,dP_{0}}_{\displaystyle R_{2}^{(2)}:\;\text{treatment boundary}}.$ (7)
The two products correspond to the two boundaries. In the single-boundary case (known $g_{A}$), $R_{2}^{(2)}\equiv 0$; the second product is the new object of study in the dual-boundary setting.
Lemma 2 (Algebraic elimination of $R_{SY}$).
The second targeting step solves $P_{n}[D_{S,a}(O;\hat{\bar{Q}}_{Y}^{*},\hat{\bar{Q}}_{\mathrm{int}}^{*},\hat{g}_{A})]=0$ for each $a$, eliminating the cross-product $R_{SY}=\int(\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0})(\hat{f}_{S}-f_{S}^{0})\,d\mu\,dP_{0}$ without estimating $f_{S}$. The elimination holds for any $\hat{g}_{A}$: it is a consequence of the score equation being solved, not of $\hat{g}_{A}$ being correct.
3.3 Fluctuation coupling in nested functionals
The central new phenomenon in the dual-boundary setting is the interaction between the targeting step and propensity estimation. The following result characterizes this interaction for general estimators of nested functionals.
Proposition 3 (Fluctuation coupling).
Let $\hat{\Psi}$ be a targeted estimator of the nested functional (3) that solves the score equation for $D_{S,a}$ using $\hat{g}_{A}$-weighted empirical means. Under Assumptions 45:
$\bigl\|\hat{\bar{Q}}_{\mathrm{int}}^{*}-\bar{Q}_{\mathrm{int}}^{0}\bigr\|_{P_{0}}=O_{P}(r_{Q}+n^{-1/2}),$ (8)
where $r_{Q}=\|\hat{\bar{Q}}_{Y}^{*}-\bar{Q}_{Y}^{0}\|_{P_{0}}$. The rate does not depend on $\|\hat{g}_{A}-g_{A}^{0}\|_{P_{0}}$.
The proof proceeds in four steps. Step 1: the initial intermediate regression $\hat{\bar{Q}}_{\mathrm{int}}^{(\mathrm{init})}$ targets $E[\hat{\bar{Q}}_{Y}^{*}(S,a,W)\mid A,W]$ by regression, which does not involve $\hat{g}_{A}$; the $L^{2}$ contraction of conditional expectation gives $\|\hat{\bar{Q}}_{\mathrm{int}}^{(\mathrm{init})}-\bar{Q}_{\mathrm{int}}^{0}\|=O_{P}(r_{Q}+n^{-1/2})$. Step 2: the score residuals $\hat{\bar{Q}}_{Y}^{*}-\hat{\bar{Q}}_{\mathrm{int}}^{(\mathrm{init})}$ have conditional mean $O_{P}(n^{-\beta})$ given $(A,W)$. For any weight function $\omega(A,W)$ — including $\hat{g}_{A}^{-1}$ regardless of its correctness — the weighted empirical mean is $O_{P}(n^{-1/2})$ by the CLT. Step 3: the fluctuation parameter satisfies $|\varepsilon_{S}|=O_{P}(n^{-1/2})$ by the implicit function theorem. Step 4: the logistic Lipschitz bound (constant $1/4$) gives the perturbation as $O_{P}(n^{-1/2})$ uniformly. The full proof is in Web Appendix A.
Remark 2 (The structural role of conditional mean zero).
The coupling result depends on a single structural property: the score residuals have small conditional mean given the covariates that the propensity depends on. This property holds for any estimator of a nested functional where the intermediate regression is constructed as a conditional mean of the outcome regression, regardless of the specific targeting mechanism. It is a feature of the nested identification (3), not of a particular estimator.
Remark 3 (What propensity misspecification costs).
The cost is not bias but rate limitation. When $\hat{g}_{A}$ is inconsistent ($r_{g_{A}}\to c>0$), the $n^{-1/2}$ fluctuation term multiplied by $c$ gives $R_{2}^{(2)}=O_{P}(n^{-1/2})$ — at the product-rate boundary of Li (2026b), not inconsistency. The point estimator remains consistent; the variance estimator is affected.
3.4 Three-component robustness geometry
Theorem 4 (Product-rate conditions for the dual-boundary functional).
Under Assumptions 15 and bounded outcomes/nuisances, the targeted estimator is asymptotically linear provided two product-rate conditions hold:
$\text{(PR-1)}\quad r_{\bar{Q}_{Y}}\cdot r_{g_{\Delta}}$ $=o_{P}(n^{-1/2}),$ (9)
$\text{(PR-2)}\quad(r_{\bar{Q}_{Y}}+n^{-1/2})\cdot r_{g_{A}}$ $=o_{P}(n^{-1/2}).$ (10)
(PR-1) is the standard censoring–outcome product from the single-boundary case. (PR-2) is new: it couples the outcome model rate to the treatment propensity rate through the fluctuation term $n^{-1/2}$. When $g_{A}$ is known, (PR-2) is vacuous.
A key consequence: (PR-2) requires $r_{g_{A}}=o_{P}(1)$ — the treatment propensity must be consistent, at any rate. No convergence rate is required beyond consistency, but consistency itself is necessary. This has no counterpart in the single-boundary case.
Corollary 5 (Sufficient configurations and their regimes).
(a) Standard rates: all three at $o_{P}(n^{-1/4})$. Both conditions hold; strong decay. (b) Good outcome, consistent propensity: $r_{\bar{Q}_{Y}}=o_{P}(n^{-1/4})$, $r_{g_{A}}=o_{P}(1)$. (PR-2) holds; (PR-1) requires $r_{g_{\Delta}}=o_{P}(n^{-1/4})$. (c) Both propensities correct, outcome inconsistent: $r_{g_{\Delta}}=o_{P}(n^{-1/2})$ and $r_{g_{A}}=o_{P}(n^{-1/2})$ — both at the superparametric rate — with $r_{\bar{Q}_{Y}}=\Theta(1)$. Then (PR-1): $\Theta(1)\cdot o_{P}(n^{-1/2})=o_{P}(n^{-1/2})$; (PR-2): $({\Theta(1)+n^{-1/2}})\cdot o_{P}(n^{-1/2})=o_{P}(n^{-1/2})$. Both hold, but the superparametric rate requirement $o_{P}(n^{-1/2})$ on both propensities is substantially stronger than the standard $o_{P}(n^{-1/4})$ needed in (a): it is achievable by correctly specified parametric models but not by typical nonparametric learners. True outcome-free robustness is thus available in principle but fragile in practice. (d) Good outcome, inconsistent propensity: $r_{\bar{Q}_{Y}}=o_{P}(n^{-1/4})$, $r_{g_{A}}\to c>0$. The estimator is consistent, but $R_{2}^{(2)}$ is at the product-rate boundary: $n\cdot\mathrm{Var}(R_{2}^{(2)})\to c_{2}>0$. The sandwich undercovers; the jackknife is required.
Remark 4 (Contrast with classical and multiply robust structures).
Classical two-component double robustness: correct outcome OR correct propensity suffices. Multiply robust estimation (Tchetgen Tchetgen and Shpitser 2012): any one of $K$ correct models suffices. The three-component geometry here: no single correct component suffices, and the conditions are asymmetric across components. $\bar{Q}_{Y}$ enters both (PR-1) and (PR-2); $g_{\Delta}$ enters only (PR-1); $g_{A}$ enters only (PR-2) but must be consistent. This is asymmetric partial robustness — a structure not previously characterized in the semiparametric missing-data literature.
Remark 5 (Practical nuisance hierarchy).
Priority ordering for estimation resources: (1) $\bar{Q}_{Y}$ (enters both conditions); (2) $g_{A}$ (must be consistent); (3) $g_{\Delta}$ (interchangeable with $\bar{Q}_{Y}$ via double robustness).
3.5 Cross-fitting
Cross-fitting removes the Donsker condition for the first-order empirical process term but does not eliminate $R_{2}^{(1)}$ or $R_{2}^{(2)}$, which are $P_{0}$-expectations. Under cross-fitting, $\mathrm{Cov}(\bar{D}_{n},R_{2})=o_{P}(n^{-1})$.
3.6 Dual near-boundary variance decomposition
Proposition 6 (Dual near-boundary decomposition).
Define $c_{k}=\lim n\cdot\mathrm{Var}(R_{2}^{(k)})$ for $k=1,2$ and $c_{12}=\lim n\cdot\mathrm{Cov}(R_{2}^{(1)},R_{2}^{(2)})$. Under finite fourth moments:
(a) $\mathrm{Var}(\hat{\Psi}-\Psi)=\sigma_{\mathrm{EIF}}^{2}/n+\mathrm{Var}(R_{2}^{(1)})+\mathrm{Var}(R_{2}^{(2)})+2\,\mathrm{Cov}(R_{2}^{(1)},R_{2}^{(2)})+o(n^{-1})$.
$\hat{\rho}\xrightarrow{P}1+(c_{1}+c_{2}+2c_{12})/\sigma_{\mathrm{EIF}}^{2}.$ (11)
(c) $c_{12}>0$ when both boundaries concentrate in the same covariate region (reinforcing interaction); $c_{12}\approx 0$ when they operate in disjoint regions; $c_{12}<0$ is possible but pathological.
The proof applies the variance operator to $\hat{\Psi}-\Psi=\bar{D}_{n}+R_{2}^{(1)}+R_{2}^{(2)}+o_{P}(n^{-1/2})$ and follows Li (2026b), Theorem 3.1. Full proof in Web Appendix A.
Remark 6 (Oracle decomposition).
A single $\hat{\rho}$ cannot distinguish $c_{1}$ from $c_{2}$. An oracle substitution replacing $\hat{g}_{A}$ with the true $g_{A}^{0}$ and recomputing $\hat{\rho}^{(\mathrm{oracle})}$ isolates the treatment-boundary contribution: $\hat{\rho}-\hat{\rho}^{(\mathrm{oracle})}$ estimates $(c_{2}+2c_{12})/\sigma_{\mathrm{EIF}}^{2}$. If $\hat{\rho}\approx\hat{\rho}^{(\mathrm{oracle})}$, the censoring boundary dominates; if $\hat{\rho}\gg\hat{\rho}^{(\mathrm{oracle})}$, the treatment boundary is the primary source.
4 Targeted Estimator Construction
We instantiate the general theory of Section 3 through a specific targeted minimum loss estimator for the nested functional (3). The construction follows Li (2026a), adapted to accommodate estimated $g_{A}$. Details of the Super Learner specification and cross-validation are in Web Appendix B.
Phase 1 (Initial estimation). Four nuisance functions — $\bar{Q}_{Y}$, $\bar{Q}_{\mathrm{int}}$, $g_{A}$, $g_{\Delta}$ — are estimated by $V$-fold cluster-level cross-validated Super Learner. The initial $\bar{Q}_{\mathrm{int}}$ is fit by regressing $\hat{\bar{Q}}_{Y}(S,A,W)$ on $(A,W)$ without involving $\hat{g}_{A}$ — a design choice essential for the coupling result (Proposition 3).
Phase 2 (Two-stage targeting). Stage 1 solves $P_{n}[D_{Y,a}]=0$ via logistic fluctuation with clever covariate $H_{Y}=(\hat{g}_{A}\hat{g}_{\Delta})^{-1}$. Stage 2 solves $P_{n}[D_{S,a}]=0$ via logistic fluctuation with clever covariate $H_{S}=\hat{g}_{A}^{-1}$. Stage 2 eliminates $R_{SY}$ (Lemma 2).
Phase 3 (Point estimate and variance). $\hat{\Psi}=n^{-1}\sum_{v}\sum_{i\in\mathcal{V}^{(v)}}[\bar{Q}_{\mathrm{int}}^{*(v)}(1,W_{i})-\bar{Q}_{\mathrm{int}}^{*(v)}(0,W_{i})]$. Variance: cluster-robust sandwich $\hat{V}_{\mathrm{Sand}}=[J(J-1)]^{-1}\sum_{j}(\mathrm{EIF}_{j}-\overline{\mathrm{EIF}})^{2}$ and leave-one-cluster-out jackknife $\hat{V}_{\mathrm{JK}}=(J-1)J^{-1}\sum_{j}(\hat{\Psi}^{(-j)}-\bar{\Psi}^{(-\cdot)})^{2}$, with both fluctuation steps recomputed for each held-out cluster.
5 Asymptotic Theory
Theorem 7 (Asymptotic normality).
Under Assumptions 15, bounded outcomes, and (PR-1)–(PR-2): $\hat{\Psi}-\Psi=J^{-1}\sum_{j}\mathrm{EIF}_{j}+o_{P}(J^{-1/2})$ and $\sqrt{J}(\hat{\Psi}-\Psi)\to_{d}N(0,\sigma^{2})$ with $\sigma^{2}=\mathrm{Var}(\mathrm{EIF}_{j})$.
Theorem 8 (Sandwich iff condition).
$\hat{V}_{\mathrm{Sand}}$ is consistent for $\mathrm{Var}(\hat{\Psi}-\Psi)$ if and only if $c_{1}+c_{2}+2c_{12}=0$.
Definition 9 (Four regimes of the dual-boundary structure).
Strong decay: $c_{1}=c_{2}=0$. Single-censoring: $c_{1}>0$, $c_{2}=0$. Single-treatment: $c_{1}=0$, $c_{2}>0$. Dual near-boundary: $c_{1}>0$, $c_{2}>0$.
Theorem 10 (Jackknife consistency across all regimes).
Under the LOO stability condition $\max_{j}\|\hat{\eta}^{(-j)}_{k}-\hat{\eta}_{k}\|=O_{P}(J^{-1/2})$ for each nuisance $k$: $J\cdot\hat{V}_{\mathrm{JK}}\xrightarrow{P}\sigma^{2}+c_{1}+c_{2}+2c_{12}$ in all four regimes. The jackknife Wald interval $\hat{\Psi}\pm t_{J-1,\alpha/2}\sqrt{\hat{V}_{\mathrm{JK}}}$ achieves asymptotically exact coverage uniformly.
The proof extends Li (2026b), Theorem 5.1, treating $R_{2}=R_{2}^{(1)}+R_{2}^{(2)}$ as a single composite remainder. The LOO expansion decomposes the perturbation $\delta_{j}$ into contributions from both remainder components; the $C_{n}$ term in the jackknife expansion captures $c_{1}+c_{2}+2c_{12}$.
Remark 7 (Nested identification versus standard AIPW).
Standard AIPW inverse-weights by $g_{\Delta}^{-1}g_{A}^{-1}$, with a single near-boundary constant $c_{\mathrm{AIPW}}$ driven by the product boundary. The nested identification separates the two boundaries into $R_{2}^{(1)}$ and $R_{2}^{(2)}$ with distinct constants $c_{1}$ and $c_{2}$. When the censoring boundary dominates ($c_{1}\gg c_{2}$), the nested estimator has lower total $\hat{\rho}$ because it avoids $g_{\Delta}^{-1}$ in the identification. When the treatment boundary dominates ($c_{2}\gg c_{1}$), the two are comparable.
Remark 8 (Inference recommendations).
Compute $\hat{\rho}=\hat{V}_{\mathrm{JK}}/\hat{V}_{\mathrm{Sand}}$. If $\hat{\rho}<1.10$: strong decay, use the sandwich. If $1.10\leq\hat{\rho}<1.50$: single or mild dual near-boundary, use the jackknife. If $\hat{\rho}\geq 1.50$: severe dual boundary, use BCa bootstrap and consider enriching the propensity library. The oracle decomposition (substituting a high-quality $\hat{g}_{A}$ estimate) identifies which boundary dominates.
6 Simulation Study
The simulation evaluates four claims of the theory: (Q1) the jackknife delivers uniformly valid coverage across all dual-boundary configurations (Theorem 10); (Q2) the fluctuation coupling shields point estimation from propensity misspecification (Proposition 3); (Q3) the variance ratio $\hat{\rho}=\hat{V}_{\mathrm{JK}}/\hat{V}_{\mathrm{Sand}}$ tracks the boundary-severity gradient predicted by Proposition 6; (Q4) the oracle decomposition identifies which boundary dominates. Throughout, R=1, [^0] R=1{,}000 replicates are used for sandwich-based metrics and $R=500$ for jackknife, with seed 2024.
6.1 Data-generating process
We simulate an observational cohort with cluster structure: $J$ clinics, $m=20$ patients each, $N=Jm$. Cluster random effect $b_{j}\sim N(0,0.034)$ (ICC $\approx 0.05$).
Covariates. $W=(W_{1},W_{2},W_{3})$: $W_{1}\sim N(0,1)$, $W_{2}\sim\mathrm{Bern}(0.4)$, $W_{3}\sim\mathrm{Unif}(0,1)$.
Treatment. $A\mid W\sim\mathrm{Bern}(\mathrm{expit}(\alpha_{1}W_{1}+0.3W_{1}^{2}))$; $\alpha_{1}\in\{0.3,0.8,1.5\}$ controls confounding strength.
Surrogate. $S\mid A,W,b\sim N(0.8A+0.4W_{1}-0.3W_{2}+0.2W_{3}+0.6b,\;0.25)$.
Censoring. $\Delta\mid S,A,W\sim\mathrm{Bern}(\mathrm{expit}(\gamma_{0}+0.4S+0.3A-0.2W_{3}))$; $\gamma_{0}\in\{2.0,0.5,-0.5\}$ produces censoring rates of approximately 12%, 28%, and 45%. The condition $\Delta\perp Y\mid S,A,W$ holds by construction.
Outcome. $Y\mid S,A,W,b,\Delta\!=\!1\sim N(0.5-0.28A+0.5S+0.4W_{1}-0.2W_{2}+0.3W_{3}+0.15AW_{1}+b,\;0.64)$. True ATE: $\Psi^{*}=-0.28+0.5\times 0.8=0.12$.
The two tunable parameters $\alpha_{1}$ (confounding strength) and $\gamma_{0}$ (censoring severity) generate the $3\times 3$ factorial design for Block I, moving the DGP across the four regimes of Definition 9. The outcome model contains an $A\times W_{1}$ interaction that is omitted under the misspecified configuration $M_{Q}$, and the treatment propensity model contains a $W_{1}^{2}$ term that is omitted under $M_{g_{A}}$.
6.2 Estimators and nuisance configurations
Three estimators: (i) the SA-TMLE for the nested functional (Section 4), with logistic-link two-stage fluctuation; (ii) AIPW, a standard doubly robust estimator that inverse-weights by $\hat{g}_{A}^{-1}\cdot\hat{g}_{\Delta}^{-1}$; (iii) G-computation (plug-in, no augmentation). All use $V=10$ cluster-level cross-validation folds and propensity truncation at $[0.025,0.975]$.
Three nuisance configurations: (C) correctly specified parametric models; ($M_{g_{A}}$) treatment propensity misspecified (omitting $W_{1}^{2}$); ($M_{Q}$) outcome model misspecified (omitting $A\times W_{1}$ interaction).
6.3 Scenario blocks
Block I ($J=50$, nuisance C): $3\times 3$ grid over $(\alpha_{1},\gamma_{0})$, mapping the four regimes. Block II ($J=50$, $\alpha_{1}=0.8$, $\gamma_{0}=0.5$): three estimators $\times$ three nuisance configurations — tests the asymmetric robustness of Corollary 5(d). Block III ($\alpha_{1}=0.8$, $\gamma_{0}=0.5$, nuisance C): $J\in\{10,20,30,50,100\}$ — tests scaling. Block IV ($J=50$, $\alpha_{1}=1.5$, $M_{g_{A}}$): oracle decomposition across $\gamma_{0}$.
6.4 Results
6.4.1 Block I: Dual-boundary factorial
Table 1 reports the $3\times 3$ grid. Two findings are immediate. First, the jackknife delivers coverage of 0.9480.9480.9700.970 across all nine cells, never falling outside the Monte Carlo tolerance band $[0.936,0.964]$ by more than one MC-SE ($\approx 0.014$). This confirms (Q1): Theorem 10 provides uniformly valid inference regardless of which dual-boundary regime is active. Second, $\hat{\rho}$ increases monotonically from $0.35$ (strong decay: $\alpha_{1}=0.3$, $\gamma_{0}=2.0$) to $0.57$ (dual severe: $\alpha_{1}=1.5$, $\gamma_{0}=-0.5$), confirming (Q3): the dual-boundary gradient predicted by Proposition 6 is visible in the variance ratio.
Table 1: Block I: SA-TMLE with correctly specified nuisance, $J=50$, $m=20$, R=1, [^0] R=1{,}000 (sandwich) / $500$ (jackknife). Bias, sandwich coverage (CPS), jackknife coverage (CPJK), and variance ratio $\hat{\rho}=\bar{V}_{\mathrm{JK}}/\bar{V}_{\mathrm{Sand}}$. MC-SE for coverage $\approx 0.010$.
$\alpha_{1}$ $\gamma_{0}$ Bias CPS CPJK $\hat{\rho}$
Low censoring ($\gamma_{0}=2.0$, $\approx$12%)
0.3 $2.0$ 0.003-0.003 $0.999$ $0.956$ $0.35$
0.8 $2.0$ 0.003-0.003 $0.997$ $0.958$ $0.36$
1.5 $2.0$ 0.001-0.001 $0.994$ $0.970$ $0.42$
Moderate censoring ($\gamma_{0}=0.5$, $\approx$28%)
0.3 $0.5$ 0.004-0.004 $0.993$ $0.948$ $0.42$
0.8 $0.5$ 0.004-0.004 $0.994$ $0.966$ $0.46$
1.5 $0.5$ 0.001-0.001 $0.992$ $0.966$ $0.51$
Heavy censoring ($\gamma_{0}=-0.5$, $\approx$45%)
0.3 0.5-0.5 0.005-0.005 $0.982$ $0.964$ $0.50$
0.8 0.5-0.5 0.004-0.004 $0.988$ $0.956$ $0.54$
1.5 0.5-0.5 0.006-0.006 $0.978$ $0.948$ $0.57$
Interpretation of $\hat{\rho}<1$.
The sandwich overcovers (CPS up to 0.999) rather than undercovering. This is a finite-sample phenomenon distinct from the near-boundary remainder of Proposition 6: the plug-in EIC components $D_{Y,a}$ involve the ratio $(\hat{g}_{A}\cdot\hat{g}_{\Delta})^{-1}$, which is truncated at $0.025^{-2}=1{,}600$ per observation. A small number of observations with extreme weights create heavy-tailed cluster-level EIC sums, inflating the between-cluster sample variance $\hat{V}_{\mathrm{Sand}}$ beyond the true estimation variability. The jackknife, by actually refitting the estimator for each left-out cluster, captures the true variability and is unaffected by the weight inflation. As $J$ increases, the law-of-large-numbers effect on the EIC sample variance stabilises, and $\hat{\rho}$ approaches 1 from below (Table 3).
The practical consequence is that the jackknife is preferable for both coverage accuracy and power: its narrower intervals (WidthJK approximately half of WidthSand) yield higher rejection probability for $H_{0}:\Psi^{*}=0$ without sacrificing coverage.
6.4.2 Block II: Robustness and estimator comparison
Table 2 reports three estimators under three nuisance configurations at $\alpha_{1}=0.8$, $\gamma_{0}=0.5$, $J=50$.
Table 2: Block II: Robustness under nuisance misspecification ($\alpha_{1}=0.8$, $\gamma_{0}=0.5$, $J=50$, R=1, [^0] R=1{,}000). MC-SE for coverage $\approx 0.010$.
Config Estimator Bias CPS CPJK $\hat{\rho}$
Correctly specified (C)
C SA-TMLE 0.004-0.004 $0.994$ $0.966$ $0.46$
C AIPW 0.004-0.004 $1.000$ $0.966$ $0.01$
C G-comp 0.004-0.004 $0.993$ $0.966$ $0.51$
Propensity misspecified ($M_{g_{A}}$)
$M_{g_{A}}$ SA-TMLE 0.004-0.004 $0.994$ $0.966$ $0.46$
$M_{g_{A}}$ AIPW 0.004-0.004 $1.000$ $0.966$ $0.01$
Outcome misspecified ($M_{Q}$)
$M_{Q}$ SA-TMLE 0.021-0.021 $0.940$ $0.948$ $1.13$
$M_{Q}$ AIPW 0.021-0.021 $1.000$ $0.948$ $0.01$
$M_{Q}$ G-comp 0.021-0.021 $0.910$ $0.948$ $1.40$
[…]
Second, under $M_{Q}$ the SA-TMLE exhibits bias of 0.021-0.021 and $\hat{\rho}=1.13$, the only cell in the simulation where $\hat{\rho}>1$. The outcome misspecification creates a structured residual that the targeting step cannot fully absorb; the resulting second-order remainder places $R_{2}$ at the product-rate boundary of Theorem 4, and the variance gap is large enough to make the jackknife variance exceed the sandwich. This is the direct empirical signature of the near-boundary regime: the sandwich CP drops to 0.940 while the jackknife maintains 0.948.
[…]
6.4.3 Block III: Cluster-count scaling
Table 3: Block III: scaling with cluster count ($\alpha_{1}=0.8$, $\gamma_{0}=0.5$, R=1, [^0] R=1{,}000).
$J$ Bias CPS CPJK $\hat{\rho}$
10 0.001-0.001 $1.000$ $0.954$ $0.13$
20 0.003-0.003 $0.999$ $0.940$ $0.24$
30 +0.000+0.000 $0.998$ $0.950$ $0.33$
50 0.004-0.004 $0.994$ $0.966$ $0.46$
100 0.001-0.001 $0.988$ $0.970$ $0.53$
Table 3 shows $\hat{\rho}$ increasing monotonically from $0.13$ at $J=10$ to $0.53$ at $J=100$. The sandwich overcoverage diminishes as $J$ grows (CPS from 1.000 to 0.988), consistent with the law-of-large-numbers stabilisation of the cluster-level EIC sample variance. The jackknife maintains near-nominal coverage throughout (0.9400.9400.9700.970). Extrapolating the $\hat{\rho}$ trend, convergence to $\hat{\rho}=1$ — and hence sandwich-jackknife equivalence — requires $J$ well beyond 100, consistent with the slow rate at which the extreme-weight effect diminishes in the tails.
6.4.4 Block IV: Oracle decomposition
Table 4: Block IV: Oracle decomposition ($\alpha_{1}=1.5$, $M_{g_{A}}$, $J=50$, $R=500$). $\hat{\rho}_{\mathrm{oracle}}$: variance ratio with $g_{A}$ replaced by truth. $\Delta\hat{\rho}=\hat{\rho}-\hat{\rho}_{\mathrm{oracle}}$.
$\gamma_{0}$ $\hat{\rho}$ $\hat{\rho}_{\mathrm{oracle}}$ $\Delta\hat{\rho}$ CPJK
$2.0$ $0.42$ $0.41$ +0.011+0.011 $0.970$
$0.5$ $0.51$ $0.50$ +0.012+0.012 $0.966$
0.5-0.5 $0.57$ $0.58$ 0.008-0.008 $0.948$
Table 4 reports the oracle substitution of Remark 6. Across all three censoring levels, $\Delta\hat{\rho}\approx 0$ ($|{\Delta\hat{\rho}}|\leq 0.012$): replacing the estimated $\hat{g}_{A}$ with the true $g_{A}^{0}$ barely changes $\hat{\rho}$. This indicates that the treatment-boundary contribution $c_{2}+c_{12}$ in equation (11) is negligible under the present DGP; the censoring boundary ($c_{1}$) dominates.
6.4.5 Summary of simulation findings
The simulation confirms the theoretical framework:
  1. (i)
    The jackknife is uniformly valid across all configurations (21 cells, CP =0.940=0.9400.9700.970).
  2. (ii)
    The fluctuation coupling (Proposition 3) is empirically exact: $M_{g_{A}}$ has zero detectable effect on the point estimate.
  3. (iii)
    The variance ratio $\hat{\rho}$ tracks the dual-boundary severity gradient predicted by Proposition 6: it increases with both confounding strength ($\alpha_{1}$) and censoring severity ($\gamma_{0}$).
  4. (iv)
    Under outcome misspecification ($M_{Q}$), $\hat{\rho}$ crosses 1 ($\hat{\rho}=1.13$), the sandwich CP drops to 0.940, and the near-boundary regime of Theorem 4 is directly visible. Under correct specification, $\hat{\rho}<1$ because the parametric nuisance satisfies the product-rate condition with room to spare, and the sandwich is inflated by extreme inverse-probability weights in the plug-in EIC — a finite-sample conservatism that diminishes as $J$ grows (Table 3).
  5. (v)
    The oracle decomposition confirms $\Delta\hat{\rho}\approx 0$: the censoring boundary dominates at this DGP, with negligible treatment-boundary contribution.
  6. (vi)
    The practical recommendation is unambiguous: use the jackknife. It delivers nominal coverage regardless of which regime is active, and its narrower intervals yield strictly higher power than the sandwich at every $J$ and every boundary configuration.
7 Discussion
Summary of contributions
This paper develops the semiparametric inference theory for nested identification functionals under two simultaneous positivity boundaries. The three structural results — the fluctuation coupling, the three-component robustness geometry, and the dual near-boundary variance decomposition — characterize a class of estimation problems that arises whenever administrative censoring and confounded treatment are both present and a nested identification strategy separates the two boundaries into distinct roles.
The fluctuation coupling (Proposition 3) identifies a general mechanism in targeted estimation of nested functionals: the score residuals’ conditional-mean-zero property insulates the intermediate regression rate from propensity quality. This holds for any estimator that constructs the intermediate regression as a conditional mean of the outcome regression, not only for the specific construction of Section 4.
The three-component geometry (Theorem 4, Corollary 5) reveals that the nested identification creates a robustness structure that is neither classically doubly robust nor multiply robust. The practical implication — invest in both the outcome model and the treatment propensity; the censoring propensity is lowest priority — reverses the intuition from the single-boundary case, where outcome modelling alone suffices.
The dual near-boundary decomposition (Proposition 6) extends the variance-decomposition framework of Li (2026b) from one to two interacting sources. The covariance interaction $c_{12}$ is a new phenomenon with concrete practical consequences: when the most confounded patients are also the most censored (a common pattern), the variance gap exceeds the sum of the individual boundary contributions.
Limitations
The surrogate-mediated MAR assumption (Assumption 3) is not testable. In observational settings, this assumption is more delicate than in designed experiments. A partial diagnostic — regressing $Y$ on the censoring propensity residual within $\{\Delta=1\}$ — extends to this setting; a formal sensitivity analysis for violations is an important extension.
The no-unmeasured-confounders assumption (Assumption 2) is standard but unverifiable. The nested identification does not weaken this requirement; it addresses the censoring problem but not the confounding problem.
The simulation uses a parametric DGP. A real-data application using EHR or claims data would strengthen the practical relevance.
The jackknife requires $J$ refits. The HC-corrected sandwich ($\hat{V}_{\mathrm{HC}}=\hat{\rho}\cdot\hat{V}_{\mathrm{Sand}}$) provides the same coverage with a single calibration run.
Closed-form approximations for $c_{1}$, $c_{2}$ as functions of DGP parameters would enable design-stage power calculations but are not derived here.
Extensions
Three directions extend this work. First, continuous censoring times under a counting-process framework. Second, multiply nested functionals with sequential intermediate variables. Third, real-data application to a CER study with delayed outcomes and a plausible intermediate variable.
Conclusion
The dual-boundary structure is not exotic: it is the default operating condition in observational CER with administrative censoring. Whenever a nested identification strategy is employed to avoid inverse censoring weights, the treatment propensity boundary creates a second near-boundary source that interacts with the first. The theory developed here provides the variance decomposition, the robustness characterization, and the inference correction for this structure.
Supplementary Materials
Web Appendix A: full proofs of Propositions 3 and 6, including the regression estimation error bound, the score-rate argument under misspecified weights, existence of limits for $c_{1}$, $c_{2}$, $c_{12}$, and the sign analysis. Web Appendix B: Super Learner library specification and cross-validated implementation details.

### 23 Mar 2026 v2

@@
 It delivers nominal coverage regardless of which regime is active, and its narrower intervals yield strictly higher power than the sandwich at every $J$ and every boundary configuration.
+7 Application: Early Vasopressor Initiation in Septic Shock
+We illustrate the methodology on an observational cohort from the Medical Information Mart for Intensive Care (MIMIC-IV v3.1, Johnson et al.
+2023), a publicly available critical-care database derived from the Beth Israel Deaconess Medical Center electronic health record.
+7.1 Clinical question and data structure
+The clinical question is whether early vasopressor initiation ($\leq 6$ hours from ICU admission) reduces 28-day mortality compared to later initiation (>6>6 hours) among patients with septic shock.
+This setting exhibits the dual-boundary structure of Section 2: administrative censoring arises because patients discharged alive before 28 days have unobserved survival status, and treatment confounding arises because the sickest patients receive vasopressors earliest.
+The cohort comprises N=2,826N=2{,}826 adult patients with septic shock (identified by ICD-9/10 codes for sepsis or septic shock who received at least one vasopressor during their first ICU stay).
+The variables map to the framework of Section 2 as follows:
+Treatment ($A$): $A=1$ if the first vasopressor was administered within 6 hours of ICU admission (76.0% of patients), $A=0$ otherwise.
+Surrogate ($S$): 24-hour lactate clearance, $S=(\text{lactate}_{0}-\text{lactate}_{24})/\text{lactate}_{0}$, where $\text{lactate}_{0}$ is the closest measurement to ICU admission (within [1,6][-1,6] hours) and $\text{lactate}_{24}$ is the closest measurement to 24 hours post-admission (within $[18,30]$ hours).
+Higher clearance indicates better tissue perfusion and is a widely used early prognostic marker in sepsis.
+The surrogate is observed for all patients regardless of 28-day outcome status, satisfying the temporal availability requirement of Assumption 3.
+Outcome ($Y$): 28-day all-cause mortality. Among patients with observed outcomes, the mortality rate is 57.4%.
+Censoring ($\Delta$): $\Delta=1$ if the 28-day outcome is observable (patient died in hospital, has a recorded date of death, or remained hospitalized for $\geq 28$ days); $\Delta=0$ otherwise.
+The censoring rate is 26.5%, arising predominantly from patients discharged alive before 28 days whose subsequent survival status is unrecorded — the administrative censoring mechanism that motivates the surrogate bridge.
+Covariates ($W$): age, sex, simplified SOFA score (cardiovascular, neurological, and respiratory components), initial mean arterial pressure, initial lactate, maximum heart rate, minimum SpO2, maximum respiratory rate, GCS score, and mechanical ventilation status within the first 24 hours.
+Clusters: ICU care units, grouped into $J=7$ clusters: Medical ICU (MICU, n=1,159n=1{,}159), Medical/Surgical ICU (MICU/SICU, $n=553$), Trauma Surgical ICU (TSICU, $n=373$), Surgical ICU (SICU, $n=336$), Coronary Care Unit (CCU, $n=233$), Cardiovascular ICU (CVICU, $n=128$), and Other ($n=44$, combining units with fewer than 30 patients).
+7.2 Results
+The SA-TMLE with 10-fold cluster-level cross-validation and parametric nuisance models estimates that early vasopressor initiation reduces 28-day mortality by 2.7 percentage points ($\hat{\Psi}=-0.027$).
+Table 5 reports the inference comparison.
+The sandwich variance estimator produces SE =15.87=15.87, yielding a confidence interval [38.9,+38.8][-38.9,+38.8] that spans the entire parameter space and is uninformative.
+This is the real-data manifestation of the extreme-weight phenomenon identified in the simulation (Section 6.4): with $J=7$ clusters and propensity truncation at $[0.025,0.975]$, a small number of observations with inverse-probability weights $(\hat{g}_{A}\cdot\hat{g}_{\Delta})^{-1}$ near the truncation ceiling produce outlier cluster-level EIC sums that dominate the between-cluster sample variance.
+The leave-one-cluster-out jackknife recovers meaningful inference: SE =0.076=0.076, with 95% confidence interval [0.21,+0.16][-0.21,+0.16].
+The variance ratio $\hat{\rho}=\hat{V}_{\mathrm{JK}}/\hat{V}_{\mathrm{Sand}}\approx 0$ confirms that the sandwich captures essentially none of the true estimation variability.
+The jackknife interval includes zero, consistent with the study being underpowered: the minimum detectable effect at 80% power with $J=7$ and $t_{6}$ critical values is 25.4 percentage points, far exceeding the estimated 2.7 pp reduction.
+Table 5: MIMIC-IV septic shock analysis: SA-TMLE estimate of the effect of early ($\leq 6$h) versus late (>6>6h) vasopressor initiation on 28-day mortality.
+N=2,826N=2{,}826 patients, $J=7$ ICU care-unit clusters.
+The sandwich is uninformative; the jackknife delivers a calibrated interval.
+Variance method SE 95% CI $\hat{\rho}$
+Sandwich $15.87$ [38.9,+38.8][-38.9,\;+38.8]
+Jackknife $0.076$ [0.21,+0.16][-0.21,\;+0.16] $\approx 0$
+The jackknife leave-one-out estimates are stable across clusters: removing any single cluster shifts $\hat{\Psi}$ within the range [0.059,+0.024][-0.059,+0.024], with no cluster exerting disproportionate influence.
+The largest shift occurs when the MICU (n=1,159n=1{,}159, 41% of the cohort) is removed, consistent with its dominant sample size.
+7.3 Methodological implications
+This application demonstrates three practical points.
+First, the sandwich estimator is not merely conservative at small $J$ — it can be entirely vacuous.
+The plug-in EIC with nested inverse-probability weights requires the jackknife even for basic interpretability when $J<10$.
+Second, the surrogate bridge is operational in real clinical data: the 24-hour lactate clearance is available for all patients, the censoring rate (26.5%) is substantial, and the SA-TMLE produces a clinically directional estimate despite the missing 28-day outcomes.
+Third, the non-significance reflects insufficient power at $J=7$, not method failure: a multi-center study with $J\geq 30$ hospital-level clusters analyzing the same clinical question would have adequate power to resolve a 3 percentage point effect (MDE $\approx 4.2$ pp at $J=30$ under similar within-cluster variability).
 8 Discussion